OpenStudy (anonymous):

Market research has shown that at a price of $230, 100 number of plasma TV’s are sold and if the price is reduced to $190, then 200 number of TV’s are sold. Find the equation that represents this demand curve and state in the form of p=mx+b (Slope Intercept Form of a Linear Equation).

OpenStudy (anonymous):

p=mx+b is p the price? if yes, plug the numbers and you'll get two equations 230 = 100m + b and 190 = 200m + b solve those equations, you'll get m = -2/5 and b=270 so the equation is p = -(2/5)x + 270 to check whether this equation is right or not, plug the (100,230) or (200,190)